# WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada

### v3

@@ 4 Classification Methodology
This section describes a range of classifiers evaluated for the 13-class (12 categories + ‘Other’) classification of WildFireCan-MMD. We investigate three categories of classification methods. First, several VLMs such as LLaVA, Qwen, smolVLM, and GPT-4o-mini are employed as zero-shot classifiers. This is the right choice when no training data is available. Then we examined classical methods such as Decision Trees (DT), Gaussian Naive Bayes (GNB), K-Nearest Neighbours (KNN) and Support Vector Machines (SVM) to investigate a case where labeled data is available but computation is limited. Finally, we experimented with a fine-tuned transformer-based classifier to assess its performance in the presence of labeled data and moderate computational resources. Further information on classical methods and VLMs can be found in Appendices E and F.
@@
Architecture of custom classifier: To choose the architecture of the custom classifier, we experimented with the architectures proposed by various authors on the CrisisMMD dataset (presented in Appendix G) and evaluated them on the WildFireCAD-MMD dataset. These included CNNs and transformers for the image modality, as well as various BERT-based transformers for the text modality. After individually determining the best model for each modality (ViT for image with f1: 64.19%, and RoBERTA for text with f1: 77.98%), various fusion techniques found in literature and included in Table G.1 were tested to improve the classification. More information about these experiments, as well as an ablation study, can be found in Appendix H.
@@
The best performing classifier consisted of two parallel encoders, an image and text encoder. The image encoder is ViT333’google/vit-base-patch16-384’, and the text encoder is RoBERTa444’roberta-base’. The output CLS token of the last hidden state from each are sent through separate projection layers to reduce the dimensionality from 768 to 512. The output is passed to the transformer fusion module consisting of two encoder layers, each with d_model=512, nhead=8, dim_feedforward=2048, dropout=0.2. A final classification head projects the dimension from 512 to 256, which is followed by ReLU activation and dropout=0.2, then finally from 256 to num_classes. Figure I.1 shows the structure of this architecture.
@@
Training and Evaluation: Three-fold cross validation with stratified sampling was used to train the model and report the results (mean $\pm$ std). The random seeds from splitting the data were kept for reproducibility. The hyperparameters used were a batch size of 8 and an image size of ($384\times{384}\times{3}$). Text sequences are padded to a max length of 144. Adam optimizer is used with a learning rate of $1e^{-5}$, weight_decay=0.01, and cross-entropy loss function. The model is fully trained after 10 epochs. The best model is saved based on f-weighted-average. The model was tested on the held-out test set (938 samples).
@@ 5 Experimental Results
Table 3 shows the calculated f-scores for the VLMs, custom classifier, and baseline models per class and as an average over all classes. Appendix J shows the confusion matrices of the custom classifiers and VLMs. All metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in the dataset.
Table 3: F1-scores per-class for zero-shot VLMs (preferred when no training data available), 3-fold cross validated f-scores for the custom architecture (when training data and moderate computation is available), and 5-fold cross validated f-scores for baseline classifiers (when training data and limited computation is available).
@@
zero-shot VLMs Classifier Baseline Classifiers
Class LLaVA GPT-4o smolVLM Qwen DT GNB KNN SVM
Evacuees 17.65 37.21 25.84 12.70 88.55±3.20 49.42±6.66 44.44±4.65 44.39±4.50 38.42±3.18
General Information 19.05 20.37 00.00 21.88 82.04±3.54 32.49±4.63 38.05±8.41 32.88±6.67 39.35±10.31
Preparedness 57.14 62.50 22.58 67.69 66.02±4.10 48.04±6.96 62.55±5.53 55.69±5.91 54.51±2.33
Weather Reports 64.52 62.39 00.00 50.00 89.01±4.51 45.27±6.21 61.70±4.80 64.31±5.07 66.69±4.24
Warnings & Status Updates 34.94 48.00 00.00 42.73 93.19±2.02 43.35±1.27 45.87±3.35 59.32±2.68 56.99±2.14
Reports of Actions of Responders 26.09 69.09 00.00 63.89 81.10±2.14 49.43±2.23 51.37±4.50 54.35±6.08 61.72±5.73
Infrastructure 36.36 37.50 00.00 50.00 87.88±1.86 51.62±5.27 66.32±5.60 68.85±4.35 70.37±3.81
Political 69.09 77.86 00.00 85.25 83.73±0.29 51.58±5.78 66.59±4.43 55.02±5.48 60.53±4.27
Insurance 88.52 86.15 93.75 90.91 87.01±1.22 87.25±5.10 90.96±4.03 87.27±5.24 92.72±3.97
Advertisement 08.33 73.17 00.00 45.16 89.37±1.71 49.42±6.66 47.84±5.22 64.63±6.17 63.96±11.29
Smoke & Air Quality 72.46 75.38 45.48 65.96 84.20±6.11 57.79±3.33 71.38±1.20 74.93±1.17 76.24±2.51
Support 27.27 56.52 05.86 43.84 77.72±2.26 38.76±7.23 59.99±7.05 45.73±6.62 44.34±9.21
Other 05.22 57.58 25.36 52.46 81.29±1.00 40.44±4.48 53.68±5.21 50.58±2.45 56.33±2.27
Weighted Average 44.69 61.17 19.86 55.94 84.48±0.69 49.30±2.01 59.56±1.10 60.90±0.80 62.65±1.29
@@
The custom classifier achieved the highest weighted average F1-score (84.48%), outperforming both zero-shot vision-language models and traditional baseline classifiers across all classes. From the results in Table 3 and Appendix J, the ‘Evacuees’ class had the lowest F1-score under the custom classifier (78.55%), likely due to high variability in both imagery and associated text. On the other hand, the ‘Insurance’ class achieved the highest performance (91.86%), reflecting the strong visual and textual consistency across samples. Other high-performing classes include ‘Political’ (89.50 ‘Smoke& Air Quality’ (89.35%), and ‘Advertisement’ (89.17%), which may be more visually or textually distinct. In contrast, mid-range scores are seen in classes like ‘Support’ (77.12%) and ‘Warnings & Status Updates’ (79.31%), which often contain dense textual overlays that may not align well with the image modality.
@@ 6 Discussion
Task-specific fine-tuning is critical: Table 3 shows that classical models fine-tuned on labeled data generally outperform zero-shot VLMs, with the custom model achieving  23% higher f-score than VLMs. Despite a modest dataset, results affirm the value of targeted annotation and model tuning, echoing findings in literature presented in Appendix G.
@@
Future Work: The dataset exhibits a significant class imbalance, and future work may involve steps to address this issue. Utilizing the unlabeled data paired with a labelled dataset in a semi-supervised setup is a promising avenue for future research Sirbu et al. (2022); Sîrbu et al. (2025). Fine-tuning of VLMs is also a promising avenue which should be explored in future works Duan et al. (2024); Zhai et al. (2024).
@@ 7 Analysis on Trends in Data
Here, we present an analysis of an additional 46,279 posts from X, posted between 2018 and 2024, using the custom classifier described in Section 4. The collection methodology was the same as explained in section 3.1, where tweets were gathered using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions. This was combined with a keyword-based search, and then the data was combined, and duplicates were removed. The full set of search queries can be seen in Appendix K.
@@
The author locations were extracted from the samples with the goal of obtaining the province-level region from which each sample was posted. These could be either automatically generated or user-provided. The queries were run through OSMNX [^5] https://osmnx.readthedocs.io/en/stable/, which allowed for the automated search of cities or towns that users provided. If a user’s location was not found, or not in Canada, it was labeled as such. The distribution of posts by province is shown in Figure 2, indicating that the majority of the data originates from British Columbia and Alberta. Additional information about this data, including a word cloud of the text, the frequency distribution of the posts by word count, can be found in Appendix L.
@@
Figure 3 (left) shows the total number of posts per week sorted by year. This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country Reuters Staff (2023); CBC News (2024). Figure 3 (right) illustrates weekly trends in predicted labels during 2023. Sharp spikes in Smoke and Air Quality and Warnings & Status Updates likely correspond to major wildfire events occurring in Alberta and BC. Warnings & Status Updates and Evacuees follow similar patterns around the week of Aug 12, which corresponds to the state of emergency declared as the city of West Kelowna was evacuated. Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 Williams (2023). Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
@@ Acknowledgment
Appendix A Data Collection
Table A.1 shows the full queries used to collect the WildFireCan-MMD data from X.
Year Query
Table A.1: The queries used for collecting the data for WildFireCan-MMD.
Appendix B Annotation
Figure B.1 shows the distribution of the labels over the classes by each independent annotator.
Figure B.1: Counts of labels by each annotator
Appendix C Topic Modeling
Figure C.1 shows examples of topics extracted by the unsupervised BERTopic. We present representative images and corresponding keywords generated by BERTopic for each example. As seen in the captions of the image, some of the topics are not useful and present non-coherent concepts, while others are useful and relevant to wildfires.
@@ Appendix C Topic Modeling
Figure C.1: Examples of useful/non-useful topics from topic modeling on our dataset.
Appendix D Prompts for Zero-Shot Classification
The full prompt used to evaluate the aforementioned VLMs is shown in Table D.1.
Table D.1: System and User prompt used for VLMs.
@@ I. Insurance (mentions of insurance)
Appendix E Baseline Classifiers
To establish a baseline, several classical methods were evaluated on the dataset, namely DT, GNB, KNN, SVM. For encoding data into numerical features for these methods, (pretrained, but not finetuned) RoBERTa and ViT were used to extract feature sets for each modality. The concatenated feature sets from the last hidden state were saved. Principle Component Analysis (PCA) was used on the feature sets to reduce the dimensionality from 768 to 250 for the SVM method as it improved the results. For DT, class_weight=’balanced’ was used. For KNN, n_neighbors=1 was used. All other hyperparameters remained default from Scikit-learn [^6] https://scikit-learn.org/stable/. Five-fold cross validation was used to train and test all methods to report a confidence interval.
Appendix F VLMs as Zero-Shot Classifiers:
Four VLMs were considered for experimentation, GPT-4o-mini Achiam et al. (2023), LLaVA Liu et al. (2023), smolVLM Marafioti et al. (2025), Qwen2.5-VL Bai et al. (2025) and tested them in a zero-shot setting. Classification prompts for all three VLMs were created. With each query to the model, a list of all categories and their descriptions are given as a prompt along with a data instance (text + image) to classify. The prompt specifies that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Appendix D and are the same for all VLMs. The same hyperparameters are used for all VLMs with temperature=0.1 num_beams=1, max_new_tokens=1024. The same test set was used for all VLMs and consisted of a stratified split with random_state=8. This single split was discovered to be a good split from previous experimentation with baseline and custom classifiers, and was necessary due to the cost of running the VLM experiments multiple times.

### v4

@@ 3.2 Taxonomy Development
3.3 Data Labeling
Annotation was performed in Label Studio [^2] https://labelstud.io/, using the defined taxonomy presented in Table 2. Examples of labeled tweets are presented in Figure 1. Each image-text pair was given a single label. In cases where the label could differ depending on whether text or image was being considered, preference was given to the modality that provided greater value for that particular sample. For example, in Figure 1(c) the image might be labeled ‘Smoke & Air Quality’, but the text might be ‘Warnings & Status Updates’. In this case, ‘Warnings & Status Updates’ would be chosen as the label as the text contains more valuable information than the image alone. If neither contained detailed information about the time or location like the text does in this example, then the sample would be labeled as ‘Smoke & Air Quality’ as that is the only value that would otherwise be provided. The size of labeled classes is presented in Table 2.
Samples were annotated by three independent annotators, all authors of this paper. The annotators are male, 20-30 year old, fluent in English and reside in Canada. First, initial rounds of manual labeling on small subsets of the data gave a good understanding of what was being discussed and shown in the data, and confidence was gained in the decided classes. Over the course of labeling, samples were flagged that contained personal information. A vote between the labels from the three annotators decided the final labels. Annotation agreement scores can be seen in Table 3. A majority agreement was reached on 88.4% of the labels. Taking a vote between all annotators for the final label and comparing the result with each annotator, we can see which annotator was closest to the voted labels. From the results, annotators 1 & 2 had the best agreement scores. This is further demonstrated by Cohen’s Kappa metric between the different pairs of annotators. Fleiss’ Kappa metric however, indicates that a moderate level of agreement is achieved between the independent annotators. In cases of disagreement, the final label was chosen from the ‘expert’ annotator (Annotator 1), who had the most experience with the dataset collection and devising the taxonomy, and also had the highest agreement score with other annotators. Additional statistics can be found in Appendix A.
@@ 4 Classification Methodology
This section describes a range of classifiers evaluated for the 13-class (12 categories + ‘Other’) classification of WildFireCan-MMD. We investigate three categories of classification methods. First, classical methods such as Decision Trees (DT), Gaussian Naive Bayes (GNB), K-Nearest Neighbours (KNN) and Support Vector Machines (SVM) are used to investigate a case where labeled data is available but computation is limited. Then, several VLMs such as LLaVA, Qwen, smolVLM, and GPT-4o-mini are employed as zero-shot classifiers. This is the right choice when no training data is available. Finally, we experimented with a fine-tuned transformer-based classifier to assess its performance in the presence of labeled data and moderate computational resources. Further information on classical methods and VLMs can be found in Appendices D and E.
@@
Architecture of custom classifier: To determine which components to use in the selected architecture, we first experimented with various unimodal classifiers for each modality, as seen in Tables 4 & 5. These included CNNs and transformers for the image modality, as well as various BERT-based transformers for the text modality. After individually determining the best model for each modality (ViT for image with f1: 64.19%, and RoBERTA for text with f1: 77.98%), various fusion techniques found in the literature and included in Table F.1 were tested to improve the classification. The results of these multimodal experiments can be seen in Table 6. More information about these experiments, as well as an ablation study, can be found in Appendix G.
@@
The best performing classifier consisted of two parallel encoders, an image and text encoder. The image encoder is ViT333’google/vit-base-patch16-384’, and the text encoder is RoBERTa444’roberta-base’. The output CLS token of the last hidden state from each are sent through separate projection layers to reduce the dimensionality from 768 to 512. The output is passed to the transformer fusion module consisting of two encoder layers, each with d_model=512, nhead=8, dim_feedforward=2048, dropout=0.2. A final classification head projects the dimension from 512 to 256, which is followed by ReLU activation and dropout=0.2, then finally from 256 to num_classes. Figure H.1 shows the structure of this architecture.
@@
Training and Evaluation: Three data splits with stratified sampling were used to train the model and report the results. The random seeds from splitting the data were kept for reproducibility. The hyperparameters used were a batch size of 8 and an image size of ($384\times{384}\times{3}$). Text sequences are padded to a max length of 144. Adam optimizer is used with a learning rate of $1e^{-5}$, weight_decay=0.01, and cross-entropy loss function. The model is fully trained after 10 epochs. The best model is saved based on f-weighted-average. The model was tested on the held-out test set (938 samples).
Table 4: Text-only classification results.
Text Model Notes F1-Weighted
’bert-base-cased’ mean-pooling 0.7779
’bert-base-cased’ CLS 0.7793
"bertweet-base" mean-pooling 0.7900
"bertweet-base" CLS 0.7927
"ModernBERT-base" mean-pooling 0.7934
"ModernBERT-base" CLS 0.7862
"roberta-base" robertaforseq 0.7718
"roberta-base" mean-pooling 0.7934
"roberta-base" mean-pooling, half-frozen 0.7940
"roberta-base" mean-pooling, custom vocab 0.7209
"roberta-base" CLS 0.7798
"roberta-base" CLS, half-frozen 0.8008
Table 5: Image-only classification results.
Image Model Notes F1-Weighted
DenseNet121 ImageNet weights 0.5733
InceptionV3 ImageNet weights 0.5804
ResNet50 ImageNet weights 0.5841
VGG16 ImageNet weights 0.5993
"facebook/deit-base-patch16-384" CLS 0.6260
"google/vit-base-patch16-384" half-frozen, CLS 0.6290
"google/vit-base-patch16-384" mean-pooling 0.6404
"google/vit-base-patch16-384" CLS 0.6419
@@ 5 Experimental Results
Tables 4 and 5 show the unimodal text/image models experimented with. All experiments were conducted using the same train/test split.Experiments were done on different final feature representations by using either mean pooling of the last hidden state or the CLS token. Freezing the first half of the encoder layers was also tried as this has been shown to increase accuracy ingle2022. All encoders had a linear layer appended for classification. For text, "roberta-base" performed best when using just the CLS token for classification and freezing the first half of the layers, as shown in Table 4. Later experimentation found that the unfrozen version performed better in the final classifier. For images, "vit-base-patch16-384" performed best when using just the CLS token for classification without any freezing. Many of the fusion experiments tested, as well as some architectures from other studies can be seen in Table 6. The same train/test split was used as before for unimodal classification. The best result, belonging to the selected classifier, is shown in bold. Table 7 shows the calculated f-scores for the VLMs, selected and baseline classifiers per class and as an average over all classes. Appendix I shows the confusion matrices of these classifiers. All metrics reported in these tables use ’weighted’ F1-score to represent the class imbalance observed in the dataset.
Table 6: Multimodal classification results.
Text Model Image Model Fusion Type Notes F1
RoBERTa+BiLSTM ViT BiLSTM+BiLinear Koshy’s model Koshy2023 0.6149
CLIP CLIP "openai/clip-vit-base-patch32" 0.7072
VILT VILT "dandelin/vilt-b32-finetuned-nlvr2" 0.7638
RoBERTa ViT Transformer RoBERTa/ViT fully frozen 0.6164
RoBERTa x2 ViT Concatenative extract text from images with OCR 0.8102
RoBERTa ViT Concatenative roberta:mean, vit:cls 0.8028
@@
Table 7: F1-scores per-class for zero-shot VLMs, baseline and selected classifiers.
Baseline Classifiers Classifier zero-shot VLMs
Class DT GNB KNN SVM w/ PCA smolVLM LLaVA Qwen GPT-4o-mini
Evacuees 39.34±7.08 45.34±2.26 45.78±4.64 31.65±4.69 88.55±3.20 25.84 17.65 12.70 37.21
General Information 27.04±6.43 39.11±3.73 36.01±0.56 33.64±6.32 82.04±3.54 00.00 19.05 21.88 20.37
Preparedness 50.27±4.05 65.04±4.15 56.25±9.15 52.24±10.58 66.02±4.10 22.58 57.14 67.69 62.50
Weather Reports 51.48±6.89 61.22±3.86 63.70±3.44 65.37±2.89 89.01±4.51 00.00 64.52 50.00 62.39
Warnings & Status Updates 43.70±3.16 46.43±3.48 58.67±0.92 56.20±1.21 93.19±2.02 00.00 34.94 42.73 48.00
Reports of Actions of Responders 47.50±4.05 51.49±2.81 54.84±5.75 64.06±2.15 81.10±2.14 00.00 26.09 63.89 69.09
Infrastructure 55.13±8.21 67.51±6.28 69.16±6.58 66.01±8.15 87.88±1.86 00.00 36.36 50.00 37.50
Political 50.04±4.69 69.18±1.74 54.64±1.51 58.89±3.39 83.73±0.29 00.00 69.09 85.25 77.86
Insurance 86.66±1.52 91.25±1.31 90.57±3.62 92.09±1.12 87.01±1.22 93.75 88.52 90.91 86.15
Advertisement 48.80±6.41 49.56±7.07 65.92±6.43 66.11±3.23 89.37±1.71 00.00 08.33 45.16 73.17
Smoke and Air Quality 58.47±2.82 72.05±3.92 75.19±1.94 76.84±1.21 84.20±6.11 45.48 72.46 65.96 75.38
Support 47.99±10.26 60.76±4.66 47.53±6.10 43.86±6.57 77.72±2.26 05.86 27.27 43.84 56.52
Other 46.37±2.53 56.17±1.76 48.85±2.65 58.93±1.65 81.29±1.00 25.36 05.22 52.46 57.58
F1 Weighted Average 50.88±1.11 60.77±1.45 61.22±1.23 62.21±0.44 84.48±0.69 19.86 44.69 55.94 61.17
@@
The custom classifier achieved the highest weighted average F1-score (84.48%), outperforming both zero-shot vision-language models and traditional baseline classifiers across all classes. From the results in Table 7 and Appendix I, the ‘Preparedness’ class had the lowest F1-score under the custom classifier (66.02±4.10%), likely due to high variability in both imagery and associated text. On the other hand, the ‘Warnings & Status Updates’ class achieved the highest performance (93.19±2.02%), reflecting the strong visual and textual consistency across samples. Other high-performing classes include ‘Advertisement’ (89.37±1.71) ‘Weather Reports’ (89.01±4.51%), and ‘Evacuees’ (88.55±3.20%), which may be more visually or textually distinct. In contrast, mid-range scores are seen in classes like ‘Support’ (77.72±2.26%) and ‘Political’ (83.73±0.29%), which often contain dense textual overlays that may not align well with the image modality.
@@ 6 Discussion
Task-specific fine-tuning is critical: Table 7 shows that classical models fine-tuned on labeled data generally outperform zero-shot VLMs, with the custom model achieving  23% higher f-score than VLMs. Despite a modest dataset, results affirm the value of targeted annotation and model tuning, echoing findings in literature presented in Appendix F.
@@
Future Work: The dataset exhibits a significant class imbalance, and future work may involve steps to address this issue. Utilizing the unlabeled data paired with a labelled dataset in a semi-supervised setup is a promising avenue for future research sirbu2022; sirbu2025. Fine-tuning of VLMs is also a promising avenue which should be explored in future works duan2024cityllava; zhai2024fine.
@@ 7 Analysis on Trends in Data
This section describes the continuation of the analysis from the work by sherritt2025cai. Here, we present an analysis of an additional 46,279 posts from X, posted between 2018 and 2024, using the custom classifier described in Section 4. The collection methodology was the same as explained in section 3.1, where tweets were gathered using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions. This was combined with a keyword-based search, and then the data was combined, and duplicates were removed. The full set of search queries can be seen in Appendix J.
@@
The author locations were extracted from the samples with the goal of obtaining the province-level region from which each sample was posted. These could be either automatically generated or user-provided. The queries were run through OSMNX [^5] https://osmnx.readthedocs.io/en/stable/, which allowed for the automated search of cities or towns that users provided. If a user’s location was not found, or not in Canada, it was labeled as such. The distribution of posts by province is shown in Figure 2, indicating that the majority of the data originates from British Columbia and Alberta. Additional information about this data, including a word cloud of the text, the frequency distribution of the posts by word count, can be found in Appendix K.
@@
Figure 3 (left) shows the total number of posts per week sorted by year. This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country reuters2023bcbfires; cbc2024jasperfire. Figure 3 (right) illustrates weekly trends in predicted labels during 2023. Sharp spikes in Smoke and Air Quality and Warnings & Status Updates likely correspond to major wildfire events occurring in Alberta and BC. Warnings & Status Updates and Evacuees follow similar patterns around the week of Aug 12, which corresponds to the state of emergency declared as the city of West Kelowna was evacuated. Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 williams2023bcwildfires. Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
@@ Acknowledgment
Appendix A Annotation
Figure A.1 shows the distribution of the labels over the classes by each independent annotator.
Figure A.1: Counts of labels by each annotator
Appendix B Topic Modeling
Figure B.1 shows examples of topics extracted by the unsupervised BERTopic. We present representative images and corresponding keywords generated by BERTopic for each example. As seen in the captions of the image, some of the topics are not useful and present non-coherent concepts, while others are useful and relevant to wildfires.
@@ Appendix B Topic Modeling
Figure B.1: Examples of useful/non-useful topics from topic modeling on our dataset.
Appendix C Prompts for Zero-Shot Classification
The full prompt used to evaluate the aforementioned VLMs is shown in Table C.1.
Table C.1: System and User prompt used for VLMs.
@@ I. Insurance (mentions of insurance)
Appendix D Baseline Classifiers
To establish a baseline, several classical methods were evaluated on the dataset, namely DT, GNB, KNN, SVM. For encoding data into numerical features for these methods, (pretrained, but not finetuned) RoBERTa and ViT were used to extract feature sets for each modality. The concatenated feature sets from the last hidden state were saved. Principle Component Analysis (PCA) was used on the feature sets to reduce the dimensionality from 768 to 250 for the SVM method as it improved the results. For DT, class_weight=’balanced’ was used. For KNN, n_neighbors=1 was used. All other hyperparameters remained default from Scikit-learn [^6] https://scikit-learn.org/stable/. Three different random splits were used to train and test all methods to report a confidence interval.
Appendix E VLMs as Zero-Shot Classifiers:
Four VLMs were considered for experimentation, GPT-4o-mini OpenAI2023, LLaVA Liu2023, smolVLM smolvlm2025, Qwen2.5-VL qwen2025 and tested in a zero-shot setting. Classification prompts for all four VLMs were created. With each query to the model, a list of all categories and their descriptions are given as a prompt along with a data instance (text + image) to classify. The prompt specifies that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Appendix C and are the same for all VLMs. The same hyperparameters are used for all VLMs with temperature=0.1 num_beams=1, max_new_tokens=1024. The same test set was used for all VLMs and consisted of a stratified split. This single split was discovered to be a good split from previous experimentation with baseline and custom classifiers, and is one of the same three splits used for these classifiers.